OpenStudy (anonymous):

i need someone to walk me through this as well: write the point-slope form of the equation of the line through the given point with the given slope... through: (4,4)... slope= undefined

OpenStudy (aroub):

ok sure..

OpenStudy (aroub):

ok if u have an undefined slope then the line is vertical

OpenStudy (anonymous):

which means?

OpenStudy (aroub):

which means if u have a vertical line the equation looks like x= ? (something)

OpenStudy (anonymous):

so... would the slope be 0? or do you mean it could be anything?

OpenStudy (anonymous):

The slope is not 0, it's undefined. It's a vertical line

OpenStudy (aroub):

no in there is no slope its undefined..

OpenStudy (anonymous):

sorry, I'm kinda young i don't get this at all...

OpenStudy (aroub):

well let polpak explain it to u ..

OpenStudy (anonymous):

okay, thanks for helping aroub :)

OpenStudy (anonymous):

If the slope is undefined the equation will look like: x = a If the slope 0 the equation will look like: y = a

OpenStudy (anonymous):

where a is the x or y part of the point they give you.

OpenStudy (anonymous):

oh! i get it, a vertical line is on the x-axis! duh! thanks

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

It goes through the x axis

OpenStudy (anonymous):

It is parallel to the y axis

OpenStudy (anonymous):

So they gave you the point (4,4) and the slope is undefined.

OpenStudy (anonymous):

So what will the equation look like?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

um... i have y-4=m(a-4)....????

OpenStudy (anonymous):

thats probably wrong

OpenStudy (anonymous):

Nope, remember what the equation is when the slope is undefined. It's different than the usual point-slope form

OpenStudy (anonymous):

That would be right for any slope except undefined.

OpenStudy (anonymous):

4=a?

OpenStudy (anonymous):

im not getting this..

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

a is the x part of the point(4,4)

OpenStudy (anonymous):

i have the answers at the back of the packet I'm doing them in, but it says the answer is 0= x-4

OpenStudy (anonymous):

undefined slope form is: x = a where a is the x part of the point they give you.

OpenStudy (anonymous):

how does it come to that?

OpenStudy (anonymous):

That's the same answer.

OpenStudy (anonymous):

Just a dumber form.

OpenStudy (anonymous):

where did the 0 come from?

OpenStudy (anonymous):

If you take my equation, and subtract 4 from both sides..

OpenStudy (anonymous):

oh.. and why would i need to do that?

OpenStudy (anonymous):

because some people are silly and think that way looks better.

OpenStudy (anonymous):

okay, i guess i have a stupid school math department then... thanks soooooo much for your help polpak!!!!! :D

OpenStudy (anonymous):

No problem